OpenStudy (anonymous):

whats 13 over 7 = to x over 11

OpenStudy (anonymous):

\[\frac{ 13 }{ 7 } = \frac{ x }{ 11 }\]

OpenStudy (anonymous):

What you want to do here is 'cross multiply' to get rid of the fractions, \[11*13 = 7 * x\]